85.5 Plot for determining freezing point depression constant 84.5 83.5 Temperature, 82.5 Solution cooling curve 81.5 SolventYou obtained the following graphical data from two experimental runs to determine the freezing point depression constant for a pure solvent. The solution was made up by adding 3.025 g of naphthalene, C10H8, to 45.320 g of pure solvent.

1.What is the freezing point temperature of the pure solvent (in oC)?

2.What is the solution freezing point temperature (in oC)?

3.What is the freezing point depression, Tfp (in oC)?

4.What is the molar mass of naphthalene (in g/mol)?

5.How many moles of naphthalene are in the solution?

6.What is the molality of the solution (in mol/kg-solvent)?

7.What is the freezing point depression constant, kfp, of the pure solvent (in oC/m)?
Answer #1

1) The solvent temperature is 82.5 ° C

2) The temperature of the solution is 79.6 ° C.

3) The depression is calculated:

ΔT = Ts - Tsl = 82.5 - 79.6 = 2.9 ° C

4) The molar mass of naphthalene is 128.2 g / mol.

5) The moles of naphthalene are calculated:

n = m / MM = 3,025 g / 128.2 g / mol = 0.024 mol

6) The molality is calculated:

m = n / Kg = 0.024 mol / 0.04532 Kg = 0.530 m

7) The constant is calculated:

Kf = ΔT / m = 2.9 ° C / 0.53 m = 5.47 ° C / m

If you liked the answer, please rate it positively, you would help me a lot, thanks.